# Deep-Learning-Driven Prefetching for Far Memory

### v1

@@ 3.2. Prediction Task Formation
FarSight aims to reduce local memory misses by prefetching memory pages that are likely to be accessed in the near future. To reduce runtime overhead and improve the accuracy of a small DL model, we use a small vocabulary size of $K$, defaulted to 64. This means that both model input and output can only have $K$ values, implying that a future map should only have $K$ entries. Below, we explain how we achieve this small vocabulary size with our patten-addressing decoupling idea.
@@
To fit the two types of inputs into the vocabulary, we take the mod of their value to the vocabulary size, $K$. We then use a history sequence of $h$ pairs of the modulo of miss page address and PCs as the model input, as shown at the bottom of Figure 4. Even though taking a mod is inevitably a lossy process, a history sequence and two types of information still allow our model to make accurate predictions.
@@
A straightforward way to model memory miss prediction is by using their memory addresses, as used by most prior ML-based memory access prediction works (voyager, ; zhang2023dart, ; pmlr-v80-hashemi18a, ; peled2015semantic, ; peled2019neural, ). While straightforward, address-based prediction requires a huge vocabulary—$2^{3}6$ for 4KB pages in 64-bit systems. In comparison, English vocabulary used by modern Large-Language Models like GPT is only 50K to 100K in size (radford2019language, ; kaisugi2023gpt4vocab, ), beyond which prediction accuracy starts to degrade even for large models. Clearly, the huge memory address vocabulary does not meet far-memory prefetching’s accuracy demands (§2).
@@
Our solution is to label possible outcomes of memory access as ordinals. Specifically, we record a vocabulary size (i.e., $K$) of possible next memory page misses after a miss happens at page $X$. Based on the model inputs as described above, our model predicts an ordinal from 0 to $K-1$, corresponding to one of the likely next page misses. We dynamically maintain a future map) for each page $X$. Each entry in the future map represents one possible page to be accessed after the miss of page $X$, and the value of the entry is the runtime virtual memory address of the page. A null future map entry represents an outcome that has not occured at the runtime yet. We delay the detailed discussion of setting up and maintaining future maps to §3.3.
Figure 4 illustrates this idea with an example. Each page is associated with its future map of size 4 (i.e., a vocabulary size of $K=4$). For the page of addr-x, its next access can be any of the four cases in the switch clause, and the model predicts one of them based on the access history. At the state of Figure 4, only addr-y have been accessed by the application; i.e., the model has previously predicted the ordinal $1$ a fault on addr-x happened, and the subsequent access of that addr-x fault is addr-y. Similarly, on the right-hand side, pages addr-b and addr-e are tree nodes and are accessed by a Breadth-First-Search program; their next accesses are their child nodes.
@@
Vocabulary size. Naturally, a program can have fewer or more possible memory pages to access than $K$ after a page is accessed. If there are fewer possibilities (e.g., pages addr-b and addr-e only have three and two possible outcomes), the model will just not yield the remaining as a predicted value. If there are more possibilities than $K$, the model will not properly capture the less frequently occurring accesses. We will detail how we train the model so that it predicts the $K$ most frequent next accesses in §3.5.
Figure 5. FarSight Threading Model
@@
We set the configurable $K$ to 64 by default, which strikes a balance of memory overhead and prediction accuracy. Note that $K$ outcomes represent $K$ 4 KB pages, which contain a $4K$ KB address range (256 KB by default). The default value of 64 works well for all our applications for a few reasons. First, small future maps allow for hot entries to be cached at CPU L1 and L2 caches, largely reducing the prediction latency. As will be discussed in §3.3, the default 64 value already necessitates the need for swapping cold future maps to far memory when local memory is small. Second, applications with repeatable behavior usually have limited possible outcomes after one page fault. For example, pointer chasing, database B-trees, common program control flows, and sorting algorithms have one to a handful of possible memory outcomes. On the other hand, a graph with high skew could have some nodes with a large number of neighbors, but the frequency of accessing these neighbors is relatively low, and failure of prefetching them does not largely impact application performance, as shown by our PageRank results (§4.1). Third, most allocators assign addresses from a range to closely requested address allocations, resulting in most accesses being within the same memory page and $K$ pages being able to host them.
DL model architecture. We adopt the Retentive Network (RetNet) model architecture (retnet, ), which unifies the benefits of Transformer (vaswani2017attention, ) and RNN (socher2011parsing, ) to achieve O(1) inference latency and O(N) inference memory space, where N is the sequence length, while maintaining good accuracy and training speed. It achieves this goal primarily by replacing the Transformer’s softmax operation with a weighted sum of the sequence’s history context, as shown in Figure 7. By applying an exponentially decaying factor to the context, it gives more weight to the recent history (i.e., attends more to recent tokens). Even though this generic decaying method may not work with all natural language cases, it fits our usages, as program behaviors are usually influenced more by recent history than distant history. More importantly, its superior inference latency and memory consumption allow us to deploy it on each CPU core (§3.3). We feed the memory address (after mod $K$) as Q and PC mod $K$ as K and V.
3.3. Thread Model and Metadata Management
Before delving into our optimizations to the prediction method 3.4, we first discuss how we efficiently integrate model prediction, prefetching, and future-map maintenance while ensuring minimal impact to foreground application performance.
CPU-based, I/O-overlapped model prediction. As discussed in §2.3, prediction on GPU is not a viable option for far-memory prefetching. Our approach is to utilize the CPU core of each application thread to perform prediction while the thread waits for far-memory I/O operations. We observe that even with effective prefetching, on-demand network accesses (misses that are not prefetched) still occur occasionally when the local memory size is small or when applications do not exhibit good locality. When an on-demand page miss occurs during a page fault, the application thread must wait for the missed page to be fetched from main memory before proceeding. We leverage this otherwise idle CPU time to perform model prediction. Doing so avoids the latency of switching to a different CPU core and eliminates any synchronization needs across different cores, making it more scalable.
Furthermore, we maintain all model weights, application memory access history (model inputs), and hot future map entries in the core’s L1 and L2 caches. This is feasible because of the small model size and small history window we choose (totalling 20 KB of weights and metadata). As such, one model prediction takes less than 600ns, significantly faster than a network round trip of around two microseconds with RDMA.
After the model prediction returns, FarSight issues an asynchronous prefetch I/O request to far memory and immediately yields the CPU core to the application thread. Prefetched pages are placed in the Linux swap cache.
Future map management. A future map is created and associated with a memory page when the page is first accessed (and thus causing a page fault). Initially, all elements in the future map are null, implying that the system does not yet know about memory addresses for future accesses. After the model makes a prediction at the time when memory page $X$ is faulted, we look up the future map of $X$ with the index of the predicted ordinal $k$. If the $k$th element in the future map is not null, FarSight issues an asynchronous prefetch request with a far-memory address as the $k$th element’s value. If the $k$th element is null, we do not perform any prefetching, as the far-memory address has yet to be determined (this is the first-time access). When the prefetched page is being accessed, it means a page fault happens with a hit in the swap cache, and page is moved from the swap cache to for regular memory accesses. At this faulting time, we update the $k$th element’s value in page $i$’s future map with the faulting page virtual memory address.
Figure 6. Indirected and Swappable Future Maps. Application pages have one-to-one mapping to future-map roots based on their virtual page number (VPN), while the mapping from future-map roots to future maps is dynamic. On the left, the blue page (VPN 3) and its associated futuremap are swapped out. On the right, the yellow page (VPN k) and its futuremap is swapped in.
With a default $K$ value of 64 and each future map entry taking 4 bytes, each 4 KB page, including those residing in far memory, requires 256 bytes, or 6.25% of memory space. When the local memory size is small, future maps can consume significant space if all are stored in the local memory. For instance, with a local memory of 20% total memory size, future maps consume 6.25% of total memory, or 31.25% of the local memory. If this significant portion of local memory cannot be used to store application in-memory data, application performance will be largely impacted.
We propose two solutions to resolve this problem, as illustrated in Figure 6. First, we make future maps swappable at the granularity of one single future map (256B for $k=64$). When an application memory page is swapped out, we also swap out its associated future map to far memory. When an application page is swapped in at prefetching or on-demand faulting time, we also swap in its future map. Second, we add a level of indirection to support fine-grained future-map memory management—a small number of non-swappable future-map root pages. Every application page has a static mapping to a future-map root based on its virtual page number. Each future-map root stores the current virtual address of the application page’s associated future map. As future maps are swapped out, new ones take their locations by establishing new maping between roots and future maps. Doing so avoids internal fragmentation and improves the memory efficiency of future maps stored at the local memory.
3.4. Prediction Method Optimization
The previous two sections details our prediction problem formation and system implementation methods. We now describe two additional optimization methods we take to further improve FarSight’s overall performance, as shown in Figure 7.
Figure 7. FarSight’s Prediction Optimization Methods Demonstrating the use of each history window to predict $s$ misses ahead of time and predicting $f=2$ pages at a time.
@@ 3.4. Prediction Method Optimization
Multi-step lookahead. So far, we assume the model predicts the immediate next missed page. With such an approach, even if we issue a far-memory access request right after the prediction, the communication delay is likely longer than when the next miss happens, making the prefetched data arrive too late to be useful. Our solution is to predict farther into the future with a look-ahead distance to cover the communication delay to prefetch application data. Specifically, we predict and prefetch the $s$th future memory miss from the current access (i.e., $s$ is the look-ahead distance). We determine $s$ by conservatively choosing a large percentile (e.g., 95%) of profiled communication delay distribution, $d$, and the average profiled inter-arrival time between two memory accesses, $l$; $s$ is $d/l$. With this conservative setup, prefetched data could arrive before it is needed but rarely after. Furthermore, to efficiently utilizing network bandwidth, we prefetch $f$ pages at a time. To know what $f$ pages to prefetch, we leverage the autoregressive nature of the RetNet (retnet, ) model to predict $f$ future misses in a sequence at a time.
@@
Model input encoding. At the time of a miss, we use the recent history window of $h$ misses as the model input sequence. A naive way to encode the history is to treat each miss as one token and perform positional encoding of these tokens starting from position 0. This encoding works for generic sequence-to-sequence problems, as each new request to the model is treated as a different sequence. However, in our environment, most tokens (past accesses) but one overlap when we move the history window by one position ($m_{3}$ to $m_{h}$ accesses overlap between the two predictions shown in Figure 7). With the naive encoding method, we could not reuse any previously computed intermediate results as the token positions have changed in the new window, which causes recomputation performance overhead.
@@
To solve this problem and improve prediction performance, we propose a new encoding method based on rotary positional encoding (su2021roformer, ), a widely used encoding algorithm. We observe that the positional relationship between tokens is important but not the absolute starting position. Instead of always starting from 0, our encoding starts from the position where the first access in the current history window (e.g., $m_{3}$ in t2 window) was at in the previous window ($m_{3}$ at the 60-degree angle). Essentially, we turn the rotary wheel by one unit of angle at every prediction step to align the same access at the same angles. This allows us to reuse the computed context of overlapped accesses (e.g., $m_{3}$ to $m_{h}$).
3.5. Model Training
Each deployed application goes through an offline training process. We start the process by executing the application with a user-supplied sample input with fully local memory on a single server. We expect the sample input to be smaller than the actual runtime inputs and can thus run fully locally. We do not run the training sample run in a far memory setup as we expect one trained model to work across different far memory settings (e.g., different local memory sizes, different network speeds). As will be shown in §4.1, a model trained with small inputs generalizes well to different larger inputs thanks to FarSight’s decoupled representation. To adapt to different inference far-memory settings, we randomly drop out (exclude) a certain percentage (e.g., 10%) of memory accesses.
@@ 3.5. Model Training
We instrument the sample run to capture all memory accesses and then train the RetNet model with this collected trace in an offline manner. The training process uses the same vocabulary size, look-ahead distance, encoding method, and history length as introduced in §3.2 and §3.4. As the sample run executes fully locally, there is no miss or prefetching. Thus, we build a future map for each accessed memory page to track its $s$th subsequent access. As we have the oracle knowledge of the whole execution, we maintain the top $K$ most frequently accessed subsequent pages in each future map. The training target is the correct index in the future map that matches the ground truth (with full trace, we have the oracle of what page will be accessed next).
4. Evaluation Results
Figure 8. MCF Latency. Higher is better.
Figure 9. XGBoost Latency. Higher is better.
Figure 10. GAP PageRank Latency. Higher is better.
Figure 11. GAP SSSP Latency. Higher is better.
Figure 12. MCF Prefetch Hits. Higher is better.
Figure 13. XGBoost Prefetch Hits. Higher is better.
Figure 14. MCF Prefetch Hit Rate.
Figure 15. XGBoost Prefetch Hit Rate.
This section presents our evaluation results of FarSight.
Implementation. We implemented FarSight with 5.5K lines of source code in the Linux kernel. FarSight currently runs on one compute and one memory node configuration, but can apply to memory pooling with multiple memory nodes. For consistent and fast inference, we implemented the inference of RetNet with AVX512, a CPU vector instruction set available in the x86 platform. Current implementation supports a max sequence length of 64 but is adaptable to higher or lower sequence lengths given different computation latency requirements. The detailed hyper-parameters of models can be found in Table 1.
@@ 4. Evaluation Results
Environments. We evaluate FarSight on our private clusters. The compute node is running on a server equipped with a 28-core Intel Xeon Gold 5512U CPU (2.1 GHz) and 16 GB RAM. The memory node is running on a server equipped with a 16-core Intel Xeon Gold 5218 CPU (2.3 GHz) with 64 GB RAM. Both servers are connected with 100 Gbps Mellanox EDR-CX4 NIC through a 100Gbps RoCE ToR switch.
Parameter Value
Number of parameters 2240
Hidden dimension ($d_{\text{model}}$) 8
@@
Number of attention heads ($d_{\text{attn}}$) 4
Number of layers ($n_{\text{layer}}$) 2
@@
Batch size ($B$) 1024
Learning rate ($\eta$) 0.003239
Loss function Cross Entropy
Optimizer AdamW
Table 1. Model configuration and training hyper-parameters.
Baselines. We compare FarSight with two baselines: FastSwap (fastswap, ) and Hermit (hermit, ). FastSwap is a swap-based far-memory system implemented in the Linux kernel. Hermit builds on top of FastSwap and improves its swap-out procedure to avoid swap-out being the application performance bottleneck. Both systems use the Linux prefetching policy, which only follows simple and strict rules for issuing prefetches for sequential accesses. We do not have any ML-based prefetching baseline, because existing ML-based prefetching solutions are all at the micro-architecture level and do not work on a real software systems like Linux (§5).
Workloads. We evaluate FarSight and baselines with four workloads, XGBoost (chen2016xgboost, ), PageRank and shortest path in GAP benchmark suite (beamer2017gapbenchmarksuite, ), and MCF (spec2006_429mcf, ).
XGBoost is a machine learning framework for gradient boosted decision trees (GBDTs). During training, it builds an ensemble of decision trees, where each successive tree attempts to correct the errors based on a specified loss function. Because of the irregular and data-dependent nature of tree traversal, it benefits from an ML-based prefetching. We run the HIGGS dataset (higgs, ) on XGBoost for a binary classification task that consumes 7.4 GB of memory.
The GAP Benchmark Suite is a collection of graph processing benchmarks designed to evaluate the performance of graph analytics systems. In our evaluation, we select two widely-used graph algorithms, Shortest Path (SSSP) and PageRank, which are commonly used in large-scale systems. These algorithms also exhibit complex memory access patterns that are challenging for rule-based prefetchers to capture. We use the suite’s built-in graph generator to create graphs ranging from 4 million to 16 million nodes. The memory consumption of these two workloads ranges from 2.2 GB to 8.2 GB.
MCF (Minimum Cost Flow) is one of the workloads from the SPEC-CPU2006 benchmark (spec2006_429mcf, ). Due to the intricate interconnections between nodes and edges in a cost-weighted graph, MCF exhibits unpredictable and highly variable memory access patterns that challenge conventional prefetching strategies. This benchmark effectively tests our predictor’s ability to make accurate decisions under diverse and dynamic input conditions. In our evaluation, we present results using a configuration with 220 MB to 390 MB of memory consumption.
4.1. End-to-End Application Performance
We first present the end-to-end application performance and prefetching effectiveness with FarSight and the two baselines.

### v2

@@ 3.1 Prediction Task Formation
To reduce runtime overhead and improve the accuracy of a small DL model, FarSight uses a small vocabulary size of $K$, defaulted to 64. Below, we explain FarSight’s prediction process and how we achieve this small vocabulary size with our pattern-addressing decoupling idea.
@@
To fit the two types of inputs into the vocabulary, we take the mod of their value to the vocabulary size, $K$. We then use a history sequence of $h$ pairs of the modulo of miss page address and PCs as the model input, as shown at the bottom of Figure 3. Although taking a mod is a lossy process, a history sequence and two types of information allow our model to make accurate predictions.
@@
[…]
@@
Our solution is to label possible outcomes of memory access as ordinals. Specifically, we record a vocabulary size (i.e., $K$) of possible next memory page misses after a miss happens at page $X$. Based on the model inputs as described above, our model predicts an ordinal from 0 to K1K-1, corresponding to one of the likely next page misses. We dynamically maintain a future map for each page $X$ in the local memory. Each entry in the future map represents one possible page to be accessed after the miss of page $X$. When a page $Y$ is accessed after $X$ and our predicted ordinal is $k$, we fill the $k$th entry in $X$’s future map with the virtual memory address of page $Y$. A null future map entry represents an outcome that has not yet occured during runtime.
Figure 3 illustrates this idea with two example code patterns. Each page is associated with its future map of size 4 (i.e., a vocabulary size of $K=4$). As an example, addr-b is a tree node that has previously been followed by accesses of addr-c, addr-e, and addr-d. Thus, the memory addresses for memory pages containing addr-c, addr-e, and addr-d have been recorded in the future map of addr-b. When addr-b is accessed again, the model uses memory access and PC history to predict one of <0,1,2,3><0,1,2,3>. If 1 is predicted, no prefetch will be performed. Otherwise, the correponding memory page will be prefetched.
@@
Vocabulary size. Naturally, a program can have fewer or more than $K$ possible memory pages to access after a page is accessed. If there are fewer possibilities (e.g., pages addr-b and addr-e only have three and two possible outcomes), some of the future map entries will not be used, wasting local memory. If there are more possibilities than $K$, the model will not properly capture the less frequently occurring accesses.
@@
We set the configurable $K$ to 64 by default, which strikes a balance of memory overhead and prediction accuracy. Note that $K$ outcomes represent $K$ 4 KB pages, which contain a $4K$ KB address range (256 KB by default). The default value of 64 works well for all our applications for a few reasons. First, small future maps allow for hot entries to be cached at CPU L1 and L2 caches, largely reducing the prediction latency. Second, applications with repeatable behavior usually have limited possible outcomes (subsequent accesses) after one page fault. For example, pointer chasing, database B-trees, common program control flows, and sorting algorithms have one to a handful of possible memory outcomes. On the other hand, a graph with high skew could have some nodes with a large number of neighbors, but the frequency of accessing these neighbors is relatively low, and failure of prefetching them does not impact application performance much, as shown by our PageRank results (§4.1). Third, most allocators assign addresses from a range to closely requested address allocations, resulting in most accesses being within the same memory page and $K$ pages being able to host them.
DL model architecture. We adopt the Retentive Network (RetNet) model architecture Sun et al. (2023), which unifies the benefits of Transformer Vaswani et al. (2017) and RNN Socher et al. (2011). It replaces the Transformer’s softmax operation with a weighted (giving more weight to more recent history) sum of the sequence’s history context, as shown in Figure 3. We feed the memory address (after mod $K$) as Q and PC mod $K$ as K and V.
The assumption that recent tokens are more important than distant ones may not hold for natural language prediction. However, it fits memory prediction well, as program behaviors are usually influenced more by recent history than distant history.
RetNet achieves O(1) inference latency and O(N) inference memory space, where N is the sequence length, while maintaining good accuracy and training speed. Its superior inference latency and memory consumption allow us to deploy it on each CPU core. In contrast, other sequence-to-sequence models such as vanilla Transformer and RNN do not fit our latency and memory needs, since their inference either grows linearly with sequence length or involves costly trigonometric computations. State Space Models (SSMs) are another class of architectures offering constant-time inference and relying primarily on matrix operations. Importantly, RetNet, which we adopt, belongs to this family, as  Gu & Dao (2023) highlights that it is a special case of a linear SSM.
3.2 Prediction Method Optimization
Based on our basic prediction process as discussed above, we propose a set pf optimization methods to further improve FarSight’s overall performance.
@@ 3.2 Prediction Method Optimization
Look-ahead, batched prefetch. So far, we assume the model predicts the immediate next missed page. With such an approach, even if we issue a far-memory access request right after the prediction, the communication delay is likely longer than when the next miss happens, making the prefetched data arrive too late to be useful. Our solution is to predict farther into the future with a look-ahead distance to cover the communication delay to prefetch application data. Specifically, we predict and prefetch the $s$th future memory miss from the current access (i.e., $s$ is the look-ahead distance). We determine $s$ by conservatively choosing a large percentile (e.g., 95%) of profiled communication delay distribution, $d$, and the average profiled inter-arrival time between two memory accesses, $l$; $s$ is $d/l$. With this conservative setup, prefetched data could arrive before it is needed but rarely after. Furthermore, to efficiently utilizing network bandwidth, we prefetch $f$ pages at a time.
@@
Model input encoding. At the time of a miss, we use the recent history window of $h$ misses as the model input sequence. A naive way to encode the history is to treat each miss as one token and perform positional encoding of these tokens starting from position 0. This encoding works well for generic sequence-to-sequence problems, as each new request to the model is treated as a different sequence. However, in our environment, most tokens (past accesses) but one (the most recent access) overlap when we move the history window by one position ($m_{3}$ to $m_{h}$ accesses overlap between the two predictions shown in Figure 3). With the naive encoding method, we would need to recompute everything since the token positions have changed in the new window.
@@
To solve this problem and improve prediction performance, we propose a new encoding method based on rotary positional encoding Su et al. (2021). Instead of always starting from 0, our encoding starts from the position where the first access in the current history window (e.g., $m_{3}$ in t2 window) was in the previous window ($m_{3}$ at the 60-degree angle). Essentially, we turn the rotary wheel by one unit of angle at every prediction step to align the same access at the same angles. This allows us to reuse the computed context of overlapped accesses (e.g., $m_{3}$ to $m_{h}$).
CPU-based background model prediction. To minimize the impact on application performance, we try to hide FarSight’s prediction time behind foreground application execution. We perform prediction at the CPU core of each application thread while the thread waits for on-demand far-memory I/O operations. We maintain all model weights, application memory access history (model inputs), and hot future map entries in the CPU core’s L1 and L2 caches. This is feasible because of the small model size and small history window we choose (totalling 20 KB of weights and metadata). As such, one model prediction takes less than 600ns, significantly faster than a network round trip of around two microseconds with RDMA. Furthermore, we perform all prefetching asynchronously.
3.3 Model Training and Input Generalization
Each deployed application goes through an offline training process once until its usage pattern significantly shifts. This per-application training practice is acceptable in our targeted on-premise data-centers where applications are used for enterprise only. In such environments, there are typically tens to hundreds of applications per data center, and huge application input drift is rare Chen et al. (2016); Panchenko et al. (2019).
@@ 3.3 Model Training and Input Generalization
To train a model, we execute the application with a user-supplied sample input with fully local memory on a single server. We expect the sample input to be smaller than the actual runtime inputs and can thus run fully locally. As will be shown in §4.1, a model trained with small inputs generalizes well to different larger inputs thanks to FarSight’s decoupled representation. We instrument the sample run to capture memory accesses and then train the RetNet model with this collected trace in an offline manner. The training process uses the same vocabulary size, look-ahead distance, encoding method, and history length as introduced in §3.1 and §3.2. As we have the oracle knowledge of the whole execution, we maintain the top $K$ most frequently accessed subsequent pages in each future map. The training target is the correct index in the future map that matches the ground truth.
Now, let us understand why a small training input could generalize to various larger inputs during run time. FarSight makes its prediction based on “paths” of memory access “branches”, essentially how a program “branches” next based on runtime history behavior. For paths visited during training, the model is likely to correctly predict the next step if the path is visited again at run time. For example, when a program has a sequence of memory accesses with no conditional branches, any training input can capture the path. Thus, as long as the training process can cover common paths of a program, FarSight can generalize well to different inputs.
For all the applications used in our evaluation, our training time is around 30 minutes on a single A6000 GPU. The low training cost and the theoretical foundation of good input generalization make FarSight a viable solution for many on-prem data centers.
4 Evaluation Results
Figure 4: MCF
performance.
Figure 5: XGBoost performance.
Figure 6: Page rank performance.
Figure 7: Shortest path performance.
@@ 4 Evaluation Results
Implementation and environments. We implemented FarSight with 5.5K lines of source code in the Linux kernel. We use a RetNet model architecture with 2240 parameters in 2 layers. We list our model hyperparameters in Table 1. Training hyper-parameters, particularly the learning rate, were selected following the guidelines in Kaplan et al. (2020), adjusted for our model’s parameter size. We evaluate FarSight on our private clusters. The compute node is running on a server equipped with a 28-core Intel Xeon Gold 5512U CPU (2.1 GHz) and 16 GB RAM. The memory node is running on a server equipped with a 16-core Intel Xeon Gold 5218 CPU (2.3 GHz) with 64 GB RAM. Both servers are connected with 100 Gbps Mellanox EDR-CX4 NIC through a 100Gbps RoCE ToR switch.
Table 1: Model configuration and training hyper-parameters.
Model Parameter Value Training Parameter Value
Number of parameters 2240 Batch size ($B$) 1024
Hidden dimension ($d_{\text{model}}$) 8 Learning rate ($\eta$) 0.003239
@@
Number of heads ($d_{\text{attn}}$) 4 Loss function Cross Entropy
Number of layers ($n_{\text{layer}}$) 2 Optimizer AdamW
@@
Attention Mechanism MHA
Baselines. We compare FarSight with three baselines: FastSwap Amaro et al. (2020), Hermit Qiao et al. (2023), and Twilight Duong et al. (2024). FastSwap is a classic swap-based far-memory system implemented in the Linux kernel. Hermit is a SoTA far-memory system that improves FastSwap’s swap-out performance. Both systems use default Linux prefetching, which only captures and prefetches sequential memory accesses. Twilight is a closed-source SoTA ML-based CPU prefetcher that has only shown results via simulation. As a CPU prefetcher, Twilight assumes visibility into every memory access—an assumption that does not hold in an OS-level far-memory system, where page tables obscure fine-grained access information. Therefore, we built a simulator based on the Twilight paper’s algorithm, captured its prefetching trace using the simulator, and replayed the trace in the runtime far-memory system.
Workloads. We evaluate FarSight and baselines with four workloads, XGBoost Chen & Guestrin (2016), PageRank and shortest path (SSSP) in GAP benchmark suite Beamer et al. (2017), and MCF from SPEC 2006 benchmark  Henning (2006). XGBoost is a gradient boosted decision tree (GBDT) framework that trains an ensemble of trees, with each tree correcting errors defined by a loss function. We run the NYC Taxi dataset Work & Donovan (2016) on XGBoost for a classification task that consumes 40.5 GB of memory. The GAP Benchmark Suite is a collection of graph processing benchmarks designed to evaluate the performance of graph analytics systems. We Select Shortest Path (SSSP) and PageRank due to their prevalence in large-scale systems, with generated graphs of 4–16 million nodes and memory usage of 22–30.5 GB. MCF (Minimum Cost Flow) from SPEC 2006 Benchmark is a standard workload for CPU architecture evaluation. We use small input graphs ranging from 220MB to 390MB to evaluate MCF with Twilight as a baseline, because running Twilight on larger workloads takes weeks or longer.
Model Training. We trained each application’s model on a smaller input that is different from the larger testing inputs. We trained MCF with the smallest graph provided by SPEC and tested it on a graph that is 3x larger. We trained PageRank and SSSP with graphs generated from the Gapbs-provided script, which are 10x smaller than the testing graphs. We trained XGBoost on a small dataset, ”Adult Census Income”, with around 50k samples; we tested it with the NYC Taxi dataset with 73 million samples.
4.1 End-to-End Application Performance